# Neural Point-based Volumetric Avatar: Surface-guided Neural Points for Efficient and Photorealistic Volumetric Head Avatar

### v1

@@ 3.1.3 Lightweight Radiance Decoding
The radiance (i.e., color $\bm{c}_{\bm{x}}$ and density $\sigma_{\bm{x}}$) at position $\bm{x}\in\mathbb{R}^{3}$ is extracted based on (up to) its $K$ nearest neighboring points and a neural decoding MLP (see Fig. 3), inspired by prior methods [41, 46]. To increase efficiency and ensure better generalization on novel expressions, we design a novel lightweight radiance decoding process that directly aggregates the neural points and pass this “average” feature to a lightweight neural decoding network, achieving 7$\sim$ 7$\times$ 7 × speedup (compared to Point-NeRF) and better renderings. Specifically, for a given shading point, we compute a weighted average of the features and relative positions of (up to) its $K$ nearest neural points inside a sphere of radius $R$ as in Point-NeRF:
$\bm{f_{x}}=\sum_{i}\frac{w_{i}}{\sum w_{i}}\bm{f}_{i},\;\;\bm{v_{x}}=\sum_{i}\frac{w_{i}}{\sum w_{i}}\bm{v}_{i}$ (2)
where $\bm{v}_{i}\in\mathbb{R}^{6}$ is the position encoding of the displacement vector between neural point $\bm{p}_{i}$ and the shading point $\bm{x}$, $w_{i}=\frac{1}{||\bm{p_{i}}-\bm{x}||_{2}}$ is the weight that is inversely proportional to the Euclidean distance between $\bm{x}$ and $\bm{p}_{i}$. Then, two shallow MLPs are applied to decode this average feature to density $\sigma_{\bm{x}}$ and view-dependent color $\bm{c_{x}}$:
$\mathcal{F}_{d}:(\bm{f_{x}},\bm{v_{x}})\rightarrow\sigma_{x},\;\;\mathcal{F}_{c}:(\bm{f_{x}},\bm{v_{x}},\bm{d})\rightarrow\bm{c_{x}}$ (3)
@@
Positional encoding [27] is applied to every dimension of the input vector in $\mathcal{F}_{d}$ and $\mathcal{F}_{c}$. Finally, a pixel color is obtained via the integration in Eq. (1).
@@
The primary difference between our radiance decoding and previous methods [41, 2] is that we omit the per-point feature processing MLP before the feature aggregation in Eq. (2). This modification provides two benefits: it results in 7$\sim$ 7$\times$ 7 × speedup and improved generalization to unseen expressions in our dynamic modeling task (Fig. 5).
@@
Figure 4: Qualitative Comparisons with State-of-the-Art Methods. Our NPVA produces more photorealistic facial renditions on held-out test expressions compared to previous state-of-the-art methods, particularly in challenging facial regions (i.e., eyes, beard, and mouth interior). The normal expression is presented in Row 1, while the extreme expressions are shown in Rows 2. The bottom left corners of the leftmost images show the shell thickness of our NPVA. The thickness for a specified face is the variance of point-to-surface distance for the points correspond it. Red (blue) indicates larger (smaller) variance. During learning, our NPVA automatically increases capacity (i.e., thicker shell) to better model the more challenging facial regions (e.g., hair and mouth interior).
3.2 Efficient Rendering and Training
3.2.1 Patch-wise Depth-guided Volume Rendering
@@
We propose a patch-wise depth-guided sampling strategy, taking into account that the envelope of a head mesh is often not very accurate and the visible facial parts could appear on different depth levels (e.g., jaw and neck). Specifically, we define a fixed connectivity on the $256^{2}$ position map $\hat{\bm{G}}_{o}$ so that we can easily rasterize a depth map $\bm{D}_{\text{rast}}$. For a ray passing through $[p_{x},p_{y}]$, we consider a local depth patch centered on it and obtain the minimum and maximum depth values $D_{\text{min}}$ and $D_{\text{max}}$. In our implementation, we consider only nine pixels, i.e., $\{\bm{D}_{\text{rast}}(p_{x}+i,p_{y}+j)\;|\;i,j\in\{-s,0,s\}\}$, where $s$ is a hyper-parameter set to 16 on $1024\times 667$ images.
@@
$\begin{split}\bm{p}(t_{i})=&\bm{o}+t_{i}\overrightarrow{\bm{d}},\\ t_{i}\sim\mathcal{U}[d_{c}+(\frac{2(i-1)}{N}&-1)r,d_{c}+(\frac{2i}{N}-1)r]\end{split}$ (4)
@@
where $\bm{o}$ is the optical center, $\overrightarrow{\bm{d}}$ indicates the view direction, $\mathcal{U}$ represents uniform distribution, $r$ is the sampling radius and is set to 20 in our implementation, and $N$ denotes the number of sampling points.
@@
Discussion: E-NeRF [16] proposes a similar pixel-wise depth-guided sampling method, reducing the required shading points for volume rendering. However, the pixel-wise depth-guided sampling method only considers the depth of the current pixel and cannot properly handle facial parts that appear on different depth levels (e.g., beard), yielding suboptimal results like mesh-based methods (see Fig. 6).
3.2.2 GEP Training Strategy
@@
For head image rendering, artifacts usually appear on several difficult but small regions (i.e., mouth, eyes), Therefore, uniformly sampling rays covering the entire head region is inefficient. To address this issue, we propose a three-stage ray sampling strategy that consists of a Grid-based uniform sampling stage to initiate the training, an Error-based importance sampling stage to refine the challenging regions, and a Patch-based sampling stage to improve perceptual quality.
@@
Grid-Sample Stage (G-Stage) In this stage, we prioritize full coverage of the images. The image is split into equal-sized grids without overlap, and we randomly sample one ray per grid G-Stage ensures uniform sampling across all regions and generates an initial model that produces reasonable results for all regions of the image. Moreover, we keep track of the error for each grid, obtaining an error map $\bm{E}$ ($8\times 8$) for later error-based importance sampling.
@@
Error-Sample Stage (E-Stage) During this E-Stage, we adjust the sampling probabilities of the grids based on the grid error map initialized in the previous G-Stage. Specifically, the sampling probability of a grid region is proportional to its grid error, resulting in an error-based importance sampling similar to [45, 34]. As a result, we allocate more computing budget to difficult facial regions (e.g., mouth interior, hair, and eyes in "Error Sample" of Fig. 2), significantly improving image quality within the same number of training epochs (see Fig.8). Note that we maintain an error map of smaller size in this stage and dynamically update the sampling probability.
@@
Patch-Sample Stage (P-Stage) In this stage, we sample rays belonging to an image patch (instead of individual pixels) so that we can apply patch-based perceptual loss [43]. Using a perceptual loss along with per-pixel losses (e.g., $L_{2}$) can help reduce image blur, resulting in sharper images and visually better results [25, 22].
@@
Table 1: Comparisons with State-of-the-Art Methods. Compared to state-of-the-art methods, NPVA achieves great improvements on MSE and PSNR (up to 1.13 over the $2^{\text{nd}}$ best) with slightly slower rendering speed. We bold (underline) the best ($2^{\text{nd}}$ best) results.
Methods Subject 1 Subject 2 Subject 3 Inference MSE $\downarrow$ PSNR $\uparrow$ MSE $\downarrow$ PSNR $\uparrow$ MSE $\downarrow$ PSNR $\uparrow$ Time (ms) PiCA 34.50 32.75 28.83 33.53 25.49 34.07 73 DAM 28.40 33.60 23.53 34.41 23.18 34.48 172 MVP 48.59 31.26 28.25 33.62 38.82 32.24 158 NPVA (Ours) 23.70 34.38 18.16 35.54 21.95 34.72 482
@@
Table 1: Comparisons with State-of-the-Art Methods. Compared to state-of-the-art methods, NPVA achieves great improvements on MSE and PSNR (up to 1.13 over the $2^{\text{nd}}$ best) with slightly slower rendering speed. We bold (underline) the best ($2^{\text{nd}}$ best) results.
@@
Table 2: Comparison with NeRF. Our NPVA produces high-fidelity rendering results comparable to NeRF with a rather faster rendering speed (70$\sim$ 70$\times$ 70 ×), when trained both on a single frame and multiple frames.
Radius - R (mm) Decoding strategy Method Time (ms)
NPVA-full (Ours) 1024$\times$1024 R=3 lightweight Patch-Depth 23.70 482
(a.1) NPVA-256 256$\times$256 R=3 lightweight Patch-Depth 46.45 397
(a.2) NPVA-1k 1024$\times$1024 R=3 lightweight Patch-Depth 26.36 423
(a.3) NPVA-2k 2048$\times$2048 R=3 lightweight Patch-Depth 27.36 506
(b) NPVA-heavy 1024$\times$1024 R=3 Point-NeRF Patch-Depth 24.51 3129
(c.1) NPVA-noDepth 1024$\times$1024 R=3 lightweight NeRF 406.49 406
(c.2) NPVA-PixDepth 1024$\times$1024 R=3 lightweight Pixel-Depth 24.92 413
(d) NPVA-noGEP 1024$\times$1024 R=3 lightweight Patch-Depth 30.08 472
Table 3: Ablation Studies. We demonstrate the impact of different components on the results of our NPVA method (evaluated using MSE). The first row shows our full method as reference. (a) shows the importance of using an extra displacement map, which cannot be replaced by simply increasing point numbers. (b) shows that our lightweight radiance decoding can produce better rendering results on held-out expressions with a faster rendering speed (7$\sim$ 7$\times$ 7 ×). (c) shows the superiority of our patch-wise depth-guided sampling method. (d) shows that our GEP ray sampling strategy achieves great performance improvements by properly allocating the computation budget and introducing perceptual loss.
3.3 Training Losses
We have a set of images $\{I^{(i)}\},i\in\{1,2,...,N\}$, their corresponding tracked meshes $\{\mathcal{M}^{(i)}\}$ , UV position maps $\{\bm{G}^{(i)}_{o}\}$ converted from $\{\mathcal{M}^{(i)}\}$, and reconstructed depth maps $\{\bm{D}^{(i)}\}$ using Metashape software [29]. Our training losses include per-pixel photometric loss $\mathcal{L}_{\text{pho}}$, patch-based perceptual loss $\mathcal{L}_{\text{per}}$ [43], coarse mesh loss $\mathcal{L}_{m}$, two depth losses $\mathcal{L}_{d}$ and $\mathcal{L}_{\text{rd}}$, and three regularization losses $\mathcal{L}_{s}$, $\mathcal{L}_{\text{disp}}$ and $\mathcal{L}_{\text{kl}}$.
The appearance losses are defined as:
$\mathcal{L}_{pho}=\sum_{p\in\mathcal{P}}||I^{(i)}_{p}-\hat{I}^{(i)}_{p}||_{2},\;\;\mathcal{L}_{per}=\text{LPIPS}(I^{(i)}_{\mathcal{P}},\hat{I}^{(i)}_{\mathcal{P}})$ (5)
$\begin{split}&\mathcal{L}_{m}=||G_{o}-\hat{G}_{o}||_{2},\\ &\mathcal{L}_{\text{rd}}=||(D^{(i)}-{D_{\text{rast}}}^{(i)})\odot M_{D_{\text{rast}}}||_{1},\\ &\mathcal{L}_{d}=\sum_{p\in\mathcal{P}}||(D^{(i)}_{p}-\hat{D}^{(i)}_{p})\odot M_{D}||_{1}\end{split}$ (6)
where $\hat{G}_{o}$ is the decoded $256^{2}$ position map, and $D_{\text{rast}}$ is a coarse depth map rasterized using $\hat{G}_{o}$, $\hat{D}$ is our fine depth map obtained using volume rendering. Note that depth masks $M_{D}$ and $M_{D_{\text{rast}}}$ are used to only penalize those pixels whose depth errors are less than a depth threshold $\delta_{D}$ (set to 10mm) to handle outliers.
We also include three regularization losses to improve our model’s generalization ability. $\mathcal{L}_{\text{disp}}=||G_{d}\odot M_{G_{d}}||_{2}$ is a regularization term on the displacement map to constrain the final points close to the surface and prevent overfitting, where $M_{G_{d}}$ is a mask to only penalize the points whose displacement values are larger than $\delta_{\text{disp}}$ (set to 10mm). $\mathcal{L}_{s}$ is a Total Variation (TV) loss applied on the $256^{2}$ UV position map $G_{o}$ to encourage a smooth surface. $\mathcal{L}_{\text{kl}}$ is the common Kullback-Leibler (KL) divergence prior applied on the latent space in VAE training.
$\begin{split}\mathcal{L}=&\;\lambda_{\text{pho}}\mathcal{L}_{\text{pho}}+\lambda_{\text{per}}\mathcal{L}_{\text{per}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{rd}}\mathcal{L}_{\text{rd}}\\ &+\lambda_{m}\mathcal{L}_{m}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{disp}}\mathcal{L}_{\text{disp}}+\lambda_{\text{kl}}\mathcal{L}_{\text{kl}}\end{split}$ (7)
3.4 Network Structures & Implementation Details
Our network is trained in a variational auto-encoding fashion [12] following DAM and PiCA [19, 23, 21]. The encoder comprises 5 and 7 convolution layers (with the last 5 layers shared) respectively and encodes a UV position map, which is converted from a coarse tracked mesh ($\sim$ 5K vertices) and an average texture map into a latent code $\bm{z}\in\mathbb{R}^{8\times 8\times 4}$ as in PiCA [23]. The average texture map is obtained from an open-mouth expression by averaging the unwrapped textures of all camera views. Note that the tracked mesh does not contain vertices for tongue and teeth.
The decoder contains 5/7/7 convolution layers and predicts a position map, a displacement map, and a feature map, all of which are used for later radiance decoding. The position map $\bm{G}_{p}$ at $1024^{2}$ represents a coarse (i.e., less detailed) surface as it is upsampled from $\bm{G}_{o}$ at $256^{2}$, which is supervised by the input coarse mesh $\mathcal{M}$. The displacement map $\bm{G_{d}}$ at $1024^{2}$ increases details to compensate for lost geometric details of the coarse mesh $\mathcal{M}$ during estimation. The $32-D$ feature map $F$ at $1024^{2}$ contains local appearance information around the point for radiance decoding.
4 Experiments
We test on the Multiface dataset [40]. It is an open-sourced multi-view human face dataset that captures high-quality facial details from a camera array. Processed data include calibrated camera parameters, tracked meshes, and unwrapped UV texture maps ($1024\times 1024$). We obtain a depth map for each frame individually using Metashape [29] based on the provided camera parameters.
Following MVP [21], we use downsampled images ($1024\times 667$) during training. Unless stated otherwise, all our experiments are trained on a subset of expressions and tested on held-out expressions (15 randomly chose expressions and fixed), resulting in $\sim$11K frames for training and $\sim$1K frames for test. Following PiCA [23], MSE and PSNR are calculated based on image pixels under the rasterized mask for evaluation.
4.1 Comparisons with State-of-the-Art Methods
We compare with DAM, PiCA, and MVP, to demonstrate the superiority of our approach via the rendered images under novel expressions. In Tab. 2, NPVA achieves the best PSNR (up to 1.13 over the $2^{\text{nd}}$ best). Fig. 4 and Fig. 10 demonstrate that NPVA produces more realistic facial renditions, especially in challenging facial regions (e.g., eyes, hair, mouth interior). Refer to our Supp. Mat. for video comparisons.
Figure 5: Effect of lightweight radiance decoding. Our lightweight radiance decoding not only gains 7$\sim$ 7$\times$ 7 × speedup over Point-NeRF’s radiance decoding, but also produces better facial renditions for unseen expressions.
4.2 Ablation Studies

### v3

@@ 3.1.3. Lightweight Radiance Decoding
The radiance (i.e., color $\bm{c}_{\bm{x}}$ and density $\sigma_{\bm{x}}$) at position $\bm{x}\in\mathbb{R}^{3}$ is extracted based on (up to) its $K$ nearest neighboring points and a neural decoding MLP (see Fig. 3), inspired by prior methods (Xu et al., 2022; Zimny et al., 2022). To increase efficiency and ensure better generalization on novel expressions, we design a novel lightweight radiance decoding process that directly aggregates the neural points and passes this “average” feature to a lightweight neural decoding network, achieving $\sim 7\times$ speedup (compared to Point-NeRF) and better renderings. Specifically, for a given shading point, we compute a weighted average of the features and relative positions of (up to) its $K$ nearest neural points inside a sphere of radius $R$ as in Point-NeRF:
(2) $\bm{f_{x}}=\sum_{i}\frac{w_{i}}{\sum w_{i}}\bm{f}_{i},\;\;\bm{v_{x}}=\sum_{i}\frac{w_{i}}{\sum w_{i}}\bm{v}_{i}$
[…]
(3) $\mathcal{F}_{d}:(\bm{f_{x}},\bm{v_{x}})\rightarrow\sigma_{x},\;\;\mathcal{F}_{c}:(\bm{f_{x}},\bm{v_{x}},\bm{d})\rightarrow\bm{c_{x}}$
@@
Positional encoding (Mildenhall et al., 2020) is applied to every dimension of the input vector in $\mathcal{F}_{d}$ and $\mathcal{F}_{c}$. Finally, a pixel color is obtained via the integration in Eq. (1).
@@
The primary difference between our radiance decoding and previous methods (Xu et al., 2022; Bai et al., 2023) is that we omit the per-point feature processing MLP before the feature aggregation in Eq. (2). This modification provides two benefits: it results in $\sim 7\times$ speedup and improved generalization to unseen expressions in our dynamic modeling task (Fig. 5).
@@
Figure 4. Qualitative Comparisons with State-of-the-Art Methods. Our NPVA produces more photorealistic facial renditions on held-out test expressions compared to previous state-of-the-art methods, particularly in challenging facial regions (i.e., eyes, beard, and mouth interior). The normal expression is presented in Row 1, while the extreme expressions are shown in Row 2. The bottom left corners of the leftmost images show the shell thickness of our NPVA. The thickness for a specified face is the variance of point-to-surface distance for the points corresponding to it. Red (blue) indicates larger (smaller) variance. During learning, our NPVA automatically increases capacity (i.e., thicker shell) to better model the more challenging facial regions (e.g., hair and mouth interior).
3.2. Efficient Rendering and Training
3.2.1. Patch-wise Depth-guided Volume Rendering
@@
We propose a patch-wise depth-guided sampling strategy, taking into account that the envelope of a head mesh is often not very accurate and the visible facial parts could appear on different depth levels (e.g., jaw and neck). Specifically, we define a fixed connectivity on the $256^{2}$ position map $\hat{\bm{G}}_{o}$ so that we can easily rasterize a depth map $\bm{D}_{\text{rast}}$. For a ray passing through $p_{x},p_{y}]$, we consider a local depth patch centered on it and obtain the minimum and maximum depth values $D_{\text{min}}$ and $D_{\text{max}}$. In our implementation, we consider only nine pixels, i.e., $\{\bm{D}_{\text{rast}}(p_{x}+i,p_{y}+j)\;|\;i,j\in\{-s,0,s\}\}$, where $s$ is a hyper-parameter set to 16 on $1024\times 667$ images.
@@
(4) $\bm{p}(t_{i})=\bm{o}+t_{i}\overrightarrow{\bm{d}},\;\;t_{i}\sim\mathcal{U}[d_{c}+(\frac{2(i-1)}{N}-1)r,d_{c}+(\frac{2i}{N}-1)r]$
@@
where $\bm{o}$ is the optical center, $\overrightarrow{\bm{d}}$ indicates the view direction, $\mathcal{U}$ represents uniform distribution, $r$ is the sampling radius and equals 20 in our experiments, and $N$ is the number of sampling points.
Discussion:
@@
E-NeRF (Lin et al., 2022) proposes a similar pixel-wise depth-guided sampling method, reducing the required shading points for volume rendering. However, this sampling method only considers the depth of the current pixel and cannot properly handle facial parts that appear on different depth levels (e.g., beard), yielding suboptimal results like mesh-based methods (see Fig. 6).
3.2.2. GEP Training Strategy
@@
For head image rendering, artifacts usually appear on several difficult but small regions (i.e., mouth, eyes); therefore, uniformly sampling rays covering the entire head region is inefficient. To address this issue, we propose a three-stage ray sampling strategy that consists of a Grid-based uniform sampling stage to initiate the training, an Error-based importance sampling stage to refine the challenging regions, and a Patch-based sampling stage to improve perceptual quality.
Grid-Sample Stage (G-Stage)
@@
In this stage, we prioritize full coverage of the images. The image is split into equal-sized grids without overlap, and we randomly sample one ray per grid G-Stage ensures uniform sampling across all regions and generates an initial model that produces reasonable results for all regions of the image. Moreover, we keep track of the error for each grid, obtaining an error map $\bm{E}$ ($8\times 8$) for later error-based importance sampling.
Error-Sample Stage (E-Stage)
@@
During this E-Stage, we adjust the sampling probabilities of the grids based on the grid error map initialized in the previous G-Stage. Specifically, the sampling probability of a grid region is proportional to its grid error, resulting in an error-based importance sampling similar to (Zhu et al., 2022; Sucar et al., 2021). As a result, we allocate more computing budget to difficult facial regions (e.g., mouth interior, hair, and eyes in "Error Sample" of Fig. 2), significantly improving image quality within the same number of training epochs (see Fig.8a). Note that we maintain an error map of a smaller size in this stage and dynamically update the sampling probability.
Patch-Sample Stage (P-Stage)
@@
In this stage, we sample rays belonging to an image patch (instead of individual pixels) so that we can apply patch-based perceptual loss (Zhang et al., 2018). Using a perceptual loss along with per-pixel losses (e.g., $L_{2}$) can help reduce image blur, resulting in sharper images and visually better results (Menon et al., 2020; Luo et al., 2021).
Methods Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Inference
MSE $\downarrow$ LPIPS $\downarrow$ MSE $\downarrow$ LPIPS $\downarrow$ MSE $\downarrow$ LPIPS $\downarrow$ MSE $\downarrow$ LPIPS $\downarrow$ MSE $\downarrow$ LPIPS $\downarrow$ Time (ms)
PiCA 34.50 0.232 28.83 0.108 25.49 0.239 30.22 0.285 21.32 0.224 73
DAM 28.40 0.208 23.53 0.088 23.18 0.183 24.38 0.248 20.21 0.185 107
MVP 48.59 0.242 28.25 0.102 38.82 0.262 36.23 0.268 25.46 0.230 144
NPVA (Ours) 23.70 0.160 18.16 0.075 21.95 0.183 21.88 0.180 17.13 0.141 482
@@
Table 1. Comparisons with State-of-the-Art Methods. Compared to state-of-the-art methods, NPVA achieves great improvements on LPIPS and MSE (up to 5.37 lower than the $2^{\text{nd}}$ best) with slightly slower rendering speed. We bold (underline) the best ($2^{\text{nd}}$ best) results.
Radius - R (mm) Decoding strategy Method Time (ms)
NPVA-full (Ours) $\times1024$ R=3 lightweight Patch-Depth 23.70 482
(a.1) NPVA-256 $\times256$ R=3 lightweight Patch-Depth 46.45 397
(a.2) NPVA-1k $\times1024$ R=3 lightweight Patch-Depth 26.36 423
(a.3) NPVA-2k $\times2048$ R=3 lightweight Patch-Depth 27.36 506
(b) NPVA-heavy $\times1024$ R=3 Point-NeRF Patch-Depth 24.51 3129
(c.1) NPVA-noDepth $\times1024$ R=3 lightweight NeRF 406.49 406
(c.2) NPVA-PixDepth $\times1024$ R=3 lightweight Pixel-Depth 24.92 413
(d) NPVA-R4 $\times1024$ R=4 lightweight Patch-Depth 23.54 563
(e) NPVA-noGEP $\times1024$ R=3 lightweight Patch-Depth 30.08 472
Table 2. Ablation Studies. We demonstrate the impact of different components on the results of our NPVA method (evaluated using MSE). The first row shows our full method as reference. (a) shows the importance of using an extra displacement map, which cannot be replaced by simply increasing point numbers. (b) shows that our lightweight radiance decoding can produce better rendering results on held-out expressions with a faster rendering speed ($\sim 7\times$). (c) shows the superiority of our patch-wise depth-guided sampling method. (d) shows that our GEP ray sampling strategy achieves great performance improvements by properly allocating the computation budget and introducing perceptual loss.
3.3. Training Losses
We have a set of images $\{I^{(i)}\},i\in\{1,2,...,N\}$, their corresponding tracked meshes $\{\mathcal{M}^{(i)}\}$ , UV position maps $\{\bm{G}^{(i)}_{o}\}$ converted from $\{\mathcal{M}^{(i)}\}$, and reconstructed depth maps $\{\bm{D}^{(i)}\}$ using Metashape software (Over et al., 2021). Our training losses include per-pixel photometric loss $\mathcal{L}_{\text{pho}}$, patch-based perceptual loss $\mathcal{L}_{\text{per}}$ (Zhang et al., 2018), coarse mesh loss $\mathcal{L}_{m}$, two depth losses $\mathcal{L}_{d}$ and $\mathcal{L}_{\text{rd}}$, and three regularization losses $\mathcal{L}_{s}$, $\mathcal{L}_{\text{disp}}$ and $\mathcal{L}_{\text{kl}}$.
The appearance losses are defined as:
[…]
where $\hat{G}_{o}$ is the decoded $256^{2}$ position map, and $D_{\text{rast}}$ is a coarse depth map rasterized using $\hat{G}_{o}$, $\hat{D}$ is our fine depth map obtained using volume rendering. Note that depth masks $M_{D}$ and $M_{D_{\text{rast}}}$ are used to penalize only those pixels whose depth errors are less than a depth threshold $\delta_{D}$ (set to 10mm) to handle outliers.
[…]
(7) $\begin{split}\mathcal{L}=&\;\lambda_{\text{pho}}\mathcal{L}_{\text{pho}}+\lambda_{\text{per}}\mathcal{L}_{\text{per}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{\text{rd}}\mathcal{L}_{\text{rd}}\\ &+\lambda_{m}\mathcal{L}_{m}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{disp}}\mathcal{L}_{\text{disp}}+\lambda_{\text{kl}}\mathcal{L}_{\text{kl}}\end{split}$
3.4. Network Structures & Implementation Details
Our network is trained in a variational auto-encoding fashion (Kingma and Ba, 2015) following DAM and PiCA (Lombardi et al., 2018; Ma et al., 2021; Lombardi et al., 2021). The encoder comprises 5 and 7 convolution layers (with the last 5 layers shared) respectively and encodes a UV position map, which is converted from a coarse tracked mesh ($\sim$ 5K vertices) and an average texture map into a latent code $\bm{z}\in\mathbb{R}^{8\times 8\times 4}$ as in PiCA (Ma et al., 2021). The average texture map is obtained from an open-mouth expression by averaging the unwrapped textures of all camera views. Note that the tracked mesh does not contain vertices for the tongue and teeth.
The decoder contains 5/7/7 convolution layers and predicts a position map, a displacement map, and a feature map, all of which are used for later radiance decoding. The position map $\bm{G}_{p}$ at $1024^{2}$ represents a coarse (i.e., less detailed) surface as it is upsampled from $\bm{G}_{o}$ at $256^{2}$, which is supervised by the input coarse mesh $\mathcal{M}$. The displacement map $\bm{G_{d}}$ at $1024^{2}$ increases details to compensate for lost geometric details of the coarse mesh $\mathcal{M}$ during estimation. The 32D32-D feature map $F$ at $1024^{2}$ contains local appearance information around the point for radiance decoding.
4. Experiments
We test on the Multiface dataset (Wuu et al., 2022). It is an open-sourced multi-view human face dataset that captures high-quality facial details from a camera array. Processed data include calibrated camera parameters, tracked meshes, and unwrapped UV texture maps ($1024\times 1024$). We obtain a depth map for each frame individually using Metashape (Over et al., 2021) based on the provided camera parameters.
Following MVP (Lombardi et al., 2021), we use downsampled images ($1024\times 667$) during training. Unless stated otherwise, all our experiments are trained on a subset of expressions and tested on held-out expressions (15 randomly chosen expressions and fixed), resulting in $\sim$11K frames for training and $\sim$1K frames for testing. Following PiCA (Ma et al., 2021), MSE and LPIPS are calculated based on image pixels under the rasterized mask for evaluation.
4.1. Comparisons with State-of-the-Art Methods
We compare with DAM, PiCA, and MVP to demonstrate the superiority of our approach via the rendered images under novel expressions. In Tab. 1, NPVA achieves the best MSE (up to 5.37 lower than the $2^{\text{nd}}$ best). Fig. 4 and Fig. 10 demonstrate that NPVA produces more realistic facial renditions, especially in challenging facial regions (e.g., eyes, beard, mouth interior). Refer to our Supp. Mat. for video comparisons.
@@ 4.1. Comparisons with State-of-the-Art Methods
Table 3. Comparison with NeRF. Our NPVA produces high-fidelity rendering results comparable to NeRF with a rather faster rendering speed ($\sim 70\times$), when trained both on a single frame and multiple frames.
4.2. Ablation Studies